OpenStudy (anonymous):

Can anyone help me with using the method of elimination to solve the system of linear equations please? 2x+y=-1 and 4x+y=-5

OpenStudy (anonymous):

Subtract the first equation from the second one, this will eliminate y. Find x.

OpenStudy (campbell_st):

let 2x+y = -1 (1) 4x + y = -5 (2) then equation (1) - equation 2 will eliminate y -2x = 4 so x = -2 substitute into either equation to find x point is (-2, 3)